Find the sum of all numbers divided by 3 between 1 and 100.

Grade:11

1 Answers

Arun
5 years ago
Hence all the numbers will be-
3, 6, 9, ......99
Now total terms
99 = 3 + (n-1) *3
96 = (n-1) *3
n = 33
Hence 
Sum of all the terms = (33/2)(3 + 99) = 33 * 51 = 1683
 
Regards
Arun (askIITians forum expert)